OpenStudy (anonymous):

Recurrence relations: Given this formula for Catalan numbers: sigma(n, k=1)( C[sub(k-1)] C[sub(n-k)] ), C[sub(0)] = 1 = C[sub(1)], C[sub(2)] = 2, what is C[sub(3)] thru 5? For C[sub(2)] I put sigma(2, k=1)( C[sub(k-1)] C[sub(2-k)] ) = sigma(2,k=1)( 1 * 1 ) = C[sub(1)] + 1 = 1 + 1 = 2, check. Then C[sub(3)] I put sigma(3, k=1)( C[sub(0)] C[sub(3-k)] ) = sigma(3,k=1)( 1 * 2 ) = 1 + 2 + 2 = 5, check. Then C[sub(4)] doesn't check so I'm not sure if I'm even doing this right. sigma(4,k=1)( 1 * C[sub(3)] ) = sigma(4,k=1)(5) = 1 + 2 + 5 + 5 = 13. It's supposed to be 14. Where did I go wrong?

OpenStudy (loser66):

\[\sum_{k=1}^{n} C_{k-1}C_{n-k}\] \(C_0=C_1=1\), \(C_2 =2\) what is \(C_3,C_4.C_5\), right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks..

OpenStudy (loser66):

If you can't type latex, show your work by Draw box below. I can't understand what you are doing.

OpenStudy (anonymous):

can't draw in an initial post tho can you?

OpenStudy (loser66):

I think you missed the left hand side. it should be \[C_{n+1}=\sum_{k=1}^{n} C_{k-1}C_{n-k}\] right?

OpenStudy (anonymous):

Nope just \(C_{n}\)

OpenStudy (anonymous):

on a n x n grid

OpenStudy (loser66):

ok, let's do \(C_3\) \[C_3 = \sum_{k=1}^{3} C_{k-1}C_{3-k}=C_0*C_2+C_1*C_1+C_2*C_0 = 1*2+1*1+2*1=5\] And you are correct, :)

OpenStudy (anonymous):

\(C_{2} = \sum\limits_{k=1}^2 C_{k-1} C_{2-k} = \sum\limits_{k=1}^2 C_{0}C_{1} = C_{1} + 1 * 1 = 2\) checks, so does \(C_{3} = \sum\limits_{k=1}^3 C_{k-1} C_{3-k} = \sum\limits_{k=1}^3 C_{0}C_{2} = C_{1} + C_{2} + 1 * 2 = 1 + 2 + 2 = 5\) Right but then 4 does not work that way

OpenStudy (anonymous):

Like my LaTeX??

OpenStudy (loser66):

yup, do C_4

OpenStudy (anonymous):

\(C_{4} = \sum\limits_{k=1}^4 C_{k-1} C_{4-k} = \sum\limits_{k=1}^4 C_{0}C_{3} = C_{1} + C_{2} + C_{3} + 1 * 5 = 1 + 2 + 5 + 5 = 13\) see? it's 14 though

OpenStudy (anonymous):

:(

OpenStudy (loser66):

I got 14

OpenStudy (anonymous):

but how

OpenStudy (loser66):

\[C_4 = \sum_{k=1}^{4} C_{k-1}C_{4-k}\\=C_0*C_3+C_1*C_2+C_2*C_0+C_3C_0\\ = 1*5+1*2+2*1+5*1=14\]

OpenStudy (anonymous):

where'd you get all those combos? didn't do that in the previous iterations?

OpenStudy (loser66):

Look at the notation, that is "sum of". It means you replace k =1, n =4 + k =2, n=4 + k =3, n=4 + k =4 , n=4, then calculate

OpenStudy (loser66):

You have 2 terms, right? \(C_{k-1}\) and \(C_{4-k}\), so, just calculate one by one, times them together when k goes from 1 to 4. Then add the combos

OpenStudy (anonymous):

ok